James can mop a warehouse in 12 hours. One day his friend Molly helped him and it only took 4.8 hours. How long would it take Molly to do it alone?
Answer by lwsshak3(11628)

James can mop a warehouse in 12 hours. One day his friend Molly helped him and it only took 4.8 hours. How long would it take Molly to do it alone?
**
let 1/12= james' work rate
let 1/x= molly's work rate
1/4.8= work rate when James and Molly work together
..
Individual work rates=work rate when working together
1/12+1/x=1/4.8
4.8/12+4.8/x=1
.4+4.8/x=1
4.8/x=1-.4=.6
x=4.8/.6=8
Ans:
It would take Molly 8 hours to finish the job alone